# Unveil Inversion and Invariance in Flow Transformer for Versatile Image Editing

### v2

@@ 4.1 Deficiency and Relation of Euler to DDIM
Essentially, if we rescale $\mathbf{x}_{t}$ with $\frac{\mathbf{x}_{t}}{\sqrt{\alpha_{t}}}$ and rewrite Eq. 5, DDIM can also be treated as the first-order ODE solver as Euler above, and there has been extensive research on optimizing the DDIM inversion in DM. However, despite the similar formulation of Euler and DDIM inversion, the performance of Euler inversion is significantly worse than the DDIM even if the Euler inversion is evaluated on the more powerful Stable Diffusion 3.5 v.s SD 1.5 as shown in Figure 3. Such phenomena are not expected and also make optimization methods [28, 17, 3] in diffusion invalid for flow inversion.
@@ 4.2 Two-stage Flow Inversion
Our principle is to find an initial latent and inversion trajectory that is as close to the generation process as possible. This has two-fold benefits: 1. Editing friendly. The latent and trajectory fit into the prior distribution of the model so that the image can be easily manipulated by changing the text prompt. 2. Easy to preserve invariance and fidelity. If the inversion trajectory deviates from the model prior, it requires more effort such as self-attention injection to preserve the invariance and fidelity. However, injecting too much self-attention from the original image can hinder the editing. Thus, we propose a two-stage flow inversion. First, we construct a pivotal inversion close to the generating process. Second, we manually add compensation at every timestep to eliminate the left mild error.
@@
Besides, ideally, the velocity $v_{\theta}$ is $\mathbf{x}_{1}-\mathbf{x}_{0}$ according to Eq. 3 that is a stable constant. This also strengthens the motivation of fixed-point iteration, and we can use this property to average the velocities to get a more stable velocity prediction. In practice, this is also equivalent to averaging the latent series $\{\mathbf{x}_{t}^{i}\}_{i=1}^{I}$.
@@
5 Flexible Invaraince Control with AdaLN
Motivation. The invariance control aims to preserve the non-target contents in the original image while allowing the text prompt to edit target regions during the editing process. Previous methods based on the U-Net mainly utilize the attention map injection [5, 11, 47] and filtering in the frequency domain [19, 2] to adapt to various editing types. However, their results still struggle to reconcile diverse editing types, especially for the rigid and non-rigid editing types. The reason is that self-attention injections such as KV injection and frequency filtering do not directly correspond to the text and are not well aligned with text changes. Thus, such invariance preservation may counteract some text-guided changes such as layout, quantity, and large global changes.
Considering that the flow transformer has a broader and more diverse text-to-image generation ability and MM-DiT joins the text and image features in self-attention, the new architecture, and the improved generation ability motivate us to design a unified invariance control mechanism for both rigid and non-rigid editing that better aligned with the text changes and enable more editing types.
@@ 5 Flexible Invaraince Control with AdaLN
The proposed module connects the features to the changes in the text prompt and can flexibly control the invariance without hindering the editing. In practice, we build such an editing module based on the MM-DiT Stable Diffusion 3.5. We show the potential of editing in various scenarios in Figure 4. The results show that the replaced text features $\hat{M}^{t}$ in AdaLN adapt to different editing scenarios. To complete our argument, we also present the investigation of Self-Attention of MM-DiT for invariance control in Figure 5. The observations are similar to previous experiences in U-Net. Attention injection in too many timesteps may prevent some non-rigid editing effects. Please refer to Appendix for details.
Figure 5: Invariance control of self-attention with different time periods. We inject different components (Q,K,V) of self-attention with different time periods to verify the ability to preserve the unedited contents (see Appendix). Here we show injecting V value more than 20% time steps hinders the editing.
@@ 6 Experiments
In this section, we systematically validate our methods in various editing types and scenarios. We also ablate the effects of two-stage inversion and iteration steps. We provide both qualitative and quantitative results. Please also refer to the Appendix for more results.
@@ 6.2 Comparison with previous editing methods
Quantitative comparison. The quantitative results are summarized in Table 1. The comparison methods are implemented with the most suitable diffusion-based inversions for high performance. P2P and MasaCtrl are based on Direct Inversion (Dir Inv) [17]. Our method achieves excellent performance in most metrics. Concretely, the results show that P2P has a better ability to preserve the contents of the source image but this also hinders the ability to edit images. As shown in CLIP scores, the P2P CLIP scores are relatively lower. In contrast, the PnP has a higher CLIP Edited score but the other metrics such as PSNR and LIPIPS are worse. This indicates that PnP cannot preserve the invariance information. In summary, our method achieves good results in both background preservation and editing. This means that our method can make accurate editing without sacrificing either the invariance or the editing ability.
@@ 7 Limitations and Board Impact
Limitation. Tuning-free image editing is a system consisting of inversion and invariance control. One limitation is that the invariance control based on AdaLN depends on the authentic inversion resembling the generation process. If the real image is out of the model’s domain and the inversion deviates too much from the generation process, the module cannot precisely control the contents since the inversion does not fit the model’s prior distribution and text-to-image alignment is not good. Another limitation is the computational efficiency of the fixed-point iteration since this requires inferencing the transformer more times but in practice, the overhead is not large. We present the failure cases in the Appendix. We show the out-of-domain image with poor inversion. In this case, the model cannot well preserve the contents and manipulate the image as expected. The results are determined by the model’s prior, which is different from the original image.
Broad imppact. Since the flow transformer has a large prior for the image in the real world, and our method can flexibly manipulate the images with text guidance, these raise concerns about the potential for creating misleading or deceptive content. This could be used maliciously in misinformation campaigns, deepfakes, or propaganda.
@@ 8 Conclusion
We investigated tuning-free image editing based on the flow transformer. We provide systematic analysis for the inversion of rectified flow and invariance modeling based on the MM-DiT. We propose a two-stage inversion strategy to reduce the approximation error of estimating the velocity field and try to find an inversion approximating the authentic generation process to enhance the editing ability. Besides, we propose the invariance control based on the AdaLN to reconcile rigid and non-rigid editing. The AdaLN control mechanism discriminately preserves the original image features according to the text changes, enabling more diverse editing types, especially for the rigid and non-rigid editing types. Both qualitative and quantitative experiments validate the effectiveness of the proposed modules. In conclusion, the proposed framework can adapt to versatile editing types and fully leverage the large generation prior of the flow transformer for tuning-free image editing. We hope our work can inspire thinking over flexible and accurate invariance control in large-scale T2I models to take full advantage of the generation prior for downstream tasks.

### v3

@@ 4.1 Deficiency and Relation of Euler to DDIM
Essentially, if we rescale $\mathbf{x}_{t}$ with $\frac{\mathbf{x}_{t}}{\sqrt{\alpha_{t}}}$ and rewrite Eq. 5, DDIM can also be treated as the first-order ODE solver as Euler above, and there has been extensive research on optimizing the DDIM inversion in DM. However, despite the similar formulation of Euler and DDIM inversion, the performance of Euler inversion is significantly worse than the DDIM even if the Euler inversion is evaluated on the more powerful Stable Diffusion 3.5 vs SD 1.5 as shown in Figure 3. Such phenomena are not expected and also make optimization methods [25, 16, 2] in diffusion invalid for flow inversion.
@@ 4.2 Two-stage Flow Inversion
Our principle is to find an initial latent and inversion trajectory that is as close to the authentic generation process as possible, trying to avoid only overfitting the original image. This has two-fold benefits: 1. Editing friendly. The latent and trajectory fit into the prior distribution of the model and text-to-image alignment is authentic so that the image can be easily manipulated by changing the text prompt. 2. Easy to preserve invariance and fidelity. If the inversion trajectory deviates from the model prior, it requires more effort such as self-attention injection to preserve the invariance and fidelity. However, injecting too much self-attention from the original image can hinder the editing. Thus, we propose a two-stage flow inversion. First, we construct a pivotal inversion close to the generating process. Second, we manually add compensation at every timestep to eliminate the left mild error.
@@
Besides, ideally, the velocity $v_{\theta}$ is $\mathbf{x}_{1}-\mathbf{x}_{0}$ according to Eq. 3 that is a stable value. This also strengthens the motivation of fixed-point iteration, and we can use this property to average the velocities to get a more stable velocity prediction. Practically, this is also equivalent to averaging the latent series $\{\mathbf{x}_{t}^{i}\}_{i=1}^{I}$.
@@
5 Flexible Invariance Control with AdaLN
Motivation. The invariance control aims to preserve the non-target contents in the original image while allow the text prompt to flexibly edit target regions. Previous methods based on the U-Net mainly utilize the attention map injection [4, 10, 39] and filtering non-target contents in predicted noise or frequency domain [18, 1] to adapt to various editing types. However, these still struggle to reconcile diverse editing types, especially for rigid and non-rigid editing. The reason is that self-attention injections such as KV injection and frequency filtering do not directly correspond to the text and are not well aligned with text changes. Thus, such invariance preservation may counteract some text-guided changes such as layout, quantity, and large global changes. Another corresponding issue is that these mechanisms require tuning hyperparameters for different editing types, which can be sensitive and inefficient.
Considering that the flow transformer has a broader and more diverse text-to-image generation ability and MM-DiT joins the text and image features in self-attention, the new architecture, and the improved generation ability motivate us to design a unified invariance control mechanism for both rigid and non-rigid editing that better aligned with the text changes and enable diverse editing types without intricate tuning for each type.
@@ 5 Flexible Invariance Control with AdaLN
The proposed module connects the features to the changes in the text prompt and can flexibly control the invariance without hindering the editing. In practice, we build such an editing module based on the MM-DiT Stable Diffusion 3.5. We show the potential of editing in various scenarios in Figure 4. The results show that the replaced text features $\hat{M}^{t}$ in AdaLN adapt to different editing scenarios. To complete our argument, we also present the investigation of Self-Attention of MM-DiT for invariance control in Figure 5. The observations are similar to previous experiences in U-Net. Attention injection in too many timesteps may prevent some non-rigid editing effects. In contrast, we keep the injection timestep $S$ the same for all editing types, and such a unified and stable timestep does not hinder non-rigid editing. Please refer to Appendix for details.
Figure 5: Invariance control of self-attention with different time periods. We inject different components (Q,K,V) of self-attention with different time periods to verify the ability to preserve the unedited contents (Appendix). Here we show injecting V value more than 20% time steps can hinder the non-rigid editing ‘sitting to standing’.
@@ 6 Experiments
We evaluate our methods in various editing types and scenarios and further analyze the effects of two-stage inversion and iteration steps with qualitative and quantitative results. Please also see the Appendix for more results.
@@ 6.2 Comparison with previous editing methods
Quantitative comparison. The quantitative results are summarized in Table 1. The comparison methods are implemented with the most suitable diffusion-based inversions for high performance. P2P and MasaCtrl are based on Direct Inversion (Dir Inv) [16]. Our method achieves good performance in most metrics. Concretely, the results show that P2P has a better ability to preserve the contents of the source image but this also hinders the ability to edit images. As shown in CLIP scores, the P2P CLIP scores are relatively lower. In contrast, the PnP has a higher CLIP Edited score but the other metrics such as PSNR and LIPIPS are worse. This indicates that PnP cannot preserve the invariance information well. In summary, our method achieves good results in both background preservation and editing. This means that our method can make accurate editing without sacrificing either the invariance or the editing ability.
@@ 7 Limitations and Board Impact
Limitation. Tuning-free image editing is a system consisting of inversion and invariance control. One limitation is that the invariance control based on AdaLN depends on the authentic inversion resembling the generation process. If the real image is out of the model’s domain and the inversion deviates too much from the generation process, the module cannot precisely control the contents since the inversion does not fit the model’s prior distribution and text-to-image alignment is mismatched. Another limitation is the computational efficiency of the fixed-point iteration since this requires inferencing the transformer more times but in practice, the overhead is not large. We present the failure cases in the Appendix. The out-of-domain images with poor inversion cannot be well preserved and manipulated as expected. The editing results are determined by the model’s prior, which is different from the original image.
Broad impact. Our method can leverage the large-scale flow transformer to edit the images with text guidance. These raise concerns about the potential for creating misleading or deceptive content such as deepfakes.
@@ 8 Conclusion
We investigated tuning-free image editing based on the flow transformer. We provide systematic analysis for the inversion of rectified flow and invariance control based on the MM-DiT. We propose a two-stage inversion strategy to reduce the approximation error of estimating the velocity field and try to find an inversion approximating the authentic generation process to enhance the editing ability. Besides, we propose the invariance control based on the AdaLN to reconcile rigid and non-rigid editing. The AdaLN control mechanism discriminately preserves the original image features according to the text changes, enabling more diverse editing types, especially for rigid and non-rigid editing types. Both qualitative and quantitative experiments validate the proposed framework can adapt to versatile editing types and fully leverage the large generation prior of the flow transformer for tuning-free image editing. We hope our work can inspire thinking over flexible and accurate invariance control in large-scale T2I models to take full advantage of the generation prior for downstream tasks.